f(x)=e^(2x)
What is f(1/4)?

Answer by stanbon(75887)

f(x)=e^(2x)
What is f(1/4)?
--------------------
f(1/4)=e^(2(1/4))
f(1/4)=e^(1/2)
Cheers,
Stan H.